# Adaptive behavior with stable synapses

### v2

@@ Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated in machine learning and reinforcement learning to synaptic plasticity, and, in general, to changes and optimization of network parameters. However, such rapid changes are not coherent with the timescales of synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical network reconfiguration. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks. We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
1 Introduction
The study of behavioral adaptation, observed in both humans and animals, has been a longstanding subject of research. The rapidity with which behaviors can change in response to new cues challenges the conventional explanation of synaptic plasticity. This led researchers to explore mechanisms that underlie such flexible adaptations [1].
@@ 1 Introduction
In-context learning (ICL) is the capacity of a model to adapt its behavior, without weight updates, to solve tasks not encountered during training. Initially, ICL was observed in architectures tailored for few-shot learning [2] or even zero-shot learning [3]. However, the game changed when it was observed that ICL emerges naturally in large-scale transformers [4, 5]. Their exceptional capability to adapt to contextual information allowed transformer-based architectures to achieve state-of-the-art performances in many domains, such as natural language processing and image analysis[6, 7, 8]. Despite some intriguing explanations [9, 10, 11], the emergence of this property remains not completely understood. In particular, it is not clear how to associate it with in-context learning abilities observed in biological networks. The major contribution of this work is the definition of a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
Limitations of transformers architecture
Despite their success, transformer architectures present many shortcomings, primarily in memory requirements, that scale quadratically with sequence length. This limits the scalability of transformers and hinders their applicability to tasks requiring the processing of long sequences, such as full-length document analysis or video understanding. To address these problems, several efficient linearized attention models have emerged, characterized by a forward pass executed in an RNN-like manner with constant inference memory costs. Recently, deep linear RNN architectures [12, 13] have yielded notable performance improvements over transformers, particularly in long-sequence tasks. Zucchet et al. [14] explore the efficacy of these deep linear gated RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and implementing in-context supervised learning.
@@
Our investigation extends this premise, aiming to unravel how in-context learning, as observed in deep learning artificial neural networks, might also manifest in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We claim that in-context learning can be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We argue that those are a biologically plausible ingredient capable of implementing a process that is similar to the attention mechanism present in transformers. Recent findings on dendritic computational properties [15] and on the complexity of pyramidal neurons dynamics [16] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [17, 18, 19, 20]. It has has been proposed that segregation of dendritic input [18] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [20] would support backpropagation in biological neurons. In [21] authors suggest that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network. Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of temporal trajectories and reinforcement learning, involving non trivial input-output temporal relations. This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behaviour in both natural and artificial intelligence domains.
Our work generalizes and provides a biologically plausible implementation for the type of networks presented in [14]. Notably, our architecture has the same order of magnitude of trainable parameters and hidden units (see Appendix B). In our approach, in-context learning (ICL) can emerge simply by tuning the readout weights of the two involved networks. In contrast, previous implementations require biologically implausible mechanisms of temporal credit assignment for the network weights to converge correctly. We show that we can side-step this problem by dividing the architecture into two separate components and introducing an additional objective, which forces the first network to approximate the gradient of the second.
2 Methods
2.1 Dynamical adaptation replaces synaptic plasticity
Consider a generic learning task, in which the response $y$ to a state $x$ and is influenced by a set of parameters $w$. The latter are adjusted in function of an internal error $e$. This can be rewritten in a generic formulation as follows:
$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f),\\ \tau_{y}\,\dot{y}&=Y(y,w,x),\\ \tau_{w}\,\dot{w}&=W(w,x,e).\end{cases}$ (1)
where $f$ is feedback from the environment (e.g. the reward, the target behavior, …). Alternatively to the standard interpretation of $w$ as synaptic weights of a neuronal network, we consider them as the state variables of a dynamical system coupled to the dynamics of $y$, expressed by the activity of an auxiliary network. For this reason, we also refer to them as “virtual” weights.
2.2 Dynamical supervised learning for temporal trajectory
We consider the task of learning a target temporal trajectory $y^{targ}(t)$, We define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset. In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$. We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$. In this case, learning can be formulated as follows:
$\begin{cases}e&=(y^{targ}-y)\\ y&=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{x}\,e\end{cases}$ (2)
where we removed the dependence on time $t$ for simplicity. The operations required are nonlinear, and usually are naturally implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights. However, here $\bm{w}$ are not actual weights but rather dynamical variables.
We propose, as possible implementation of this, that such non-linear functions are computed by two neural networks networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\bm{x},e)\simeq W(\bm{x},e)=\bm{x}\,e\end{cases}$ (3)
In particular, we used two RNNs (see details below in the following Method subsections) that receive $\bm{x},\bm{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm). In addition, we introduce a novel concept known as a gain-modulated network. This concept is inspired by the remarkable ability observed in biological neurons, particularly L5 neurons as discussed in previous studies (e.g., [22, 21]). These neurons exhibit the capacity to non-linearly integrate segregated inputs, a process critical for various cognitive functions. More details in the section "Gain Modulated Reservoir Computing". We empirically demonstrate that this network architecture outperforms standard RNNs in approximating and generalising the required virtual update rules described above, suggesting that gain modulation might be an important requirement for the adaptive behaviour observed in biological agents. The formulation shown above can only be used to tackle linear problems, by learning linear relationships between $\bm{x}$ and $y$. However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
2.3 Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments. We start by defining a policy network, denoted as $\bm{\pi}=\mathrm{softmax}(\bm{y})$ which implements a policy mapping the agent state encoded by the vector $\bm{x}$ to a probability distribution over actions. For the sake of simplicity, We assume a linear agent such that $\bm{y}=\bm{w}\cdot\bm{x}$. This assumption is done without loss of generality since this could be easily extended by resorting to a reservoir computer as an intermediate layer as described above. The policy depends on the virtual weights $\bm{w}$, determined by the activity of an additional RNN. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behaviour in real time. This can be formalized in a formulation that is very similar to the one used above, by changing the definition of $e(t)$ as follows:
$\begin{cases}\bm{e}&=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\\ \bm{y}&=\bm{w}\cdot\bm{x}\\ \tau_{w}\,\dot{\bm{w}}&=\bm{e}\odot\bm{x}\end{cases}$ (4)
This is the dynamics obtained by evaluating the policy gradient with respect to the virtual weights $\bm{w}$. $a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones. $\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [23, 24]) at time $t$. It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero.
This formulation holds for a null discount factor [23, 24]), we refer to the Appendix section C for the description of the general case. The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}(\bm{y})\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (5)
2.4 Algorithm Distillation
To demonstrate the ability of this architecture to learn in context, we train networks Y and W using an Algorithmic Distillation protocol, as defined in [25]. Namely, we consider a family of environments $\mathcal{A}$ with different reward distributions  [^1] A family of environments is a set of possible environments with the same action-state transitions but different reward distributions.. We then train the aforementioned networks using $N_{s}$ learning histories obtained through policy gradient methods from a subset of these environments, designated as $\mathcal{A}_{\text{ID}}$. The network performance is then tested out of distribution, i.e., on the rest of the environments $\mathcal{A}_{\text{OOD}}:=\mathcal{A}\setminus\mathcal{A}_{\text{ID}}$
2.5 Gain modulated reservoir computing (GM-RC)
Reservoir Computing
In this approach, a random RNN is employed to extract features from a time-dependent signal $\bm{x}(t)$. The dynamics of the RNN describe the evolution of $N$ hidden units $\bm{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
$\tau_{z}\dot{\bm{z}}=\phi\left(J\bm{z}+R\bm{x}\right)-\bm{z}$ (6)
The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([26]). Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively. Subsequently, the features extracted by the network can be utilized to predict a target signal $\bm{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\bm{y}^{targ}(t)-\Theta\bm{z}(t)|^{2}_{2}$ is minimized. The reservoir is then implementing a map
$\bm{y}(t)=\mathrm{RNN}_{\Theta}(\{\bm{x}(s)\}_{s\leq t})$ (7)
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\bm{y}=\mathrm{RNN}_{\Theta}(\{\bm{x}\})$, dropping the temporal dependencies. 222 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [27].
When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit. In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\bm{x}\right)$. Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights. This architecture is also referred to in the literature as the Extreme Learning Machine [28].
Gain modulated network architecture
Analysing Equations (3) and (5) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters. Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC). This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [22, 29]. Here, we consider an additional input source $\bm{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$. Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope. Consequently, the resulting RNN equation is formulated as:
$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+\beta\cdot R^{\text{ap}}\bm{x}^{\text{ap}}+R\bm{x})\right)-\bm{z}$ (8)
Where $\bm{b}^{\text{ap}}$ is a constant bias vector. The hyperparameters $\alpha,\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs. Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation. Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies, we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network. The choice for the name $\bm{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [22].
3 Results
Computing resources
The experiments of this paper were executed on a Macbook pro M3 CPU 12-core with 36 GB of RAM, and on a Macbook pro 2,9 GHz 6-Core Intel Core i9 with 32 GB of RAM.

### v4

@@ Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated learning theories to synaptic plasticity. However, such rapid changes are not coherent with the timescales of biological synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical reconfiguration of the network involved, without changing its weights. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks, such as in pyramidal neurons thanks to input segregation and dendritic amplification. We propose a constructive approach to induce in-context learning in an architecture composed of recurrent networks with gain modulation, demonstrating abilities inaccessible to standard networks. In particular, we show that, such architecture can dynamically implement standard gradient-based by encoding weight changes in the activity of another network. We argue that, while these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the level of neural circuits. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
Introduction
Adaptive Behavior in Biological Agents
The study of adaptive behavior in biological agents, including humans and animals, has been a longstanding subject of research [1]. Observations of rapid behavioral shifts in response to environmental cues have raised questions about whether traditional mechanisms, such as synaptic plasticity, can fully account for such flexibility.
A broad range of neural and behavioral data indicates the existence of distinct systems for behavioral choice [2, 3]. The conventional view posits that subcortical areas support habitual, stimulus-based responses, enabling reflex-based and rapid reactions. In contrast, higher cortical areas, such as the prefrontal cortex, are associated with reflective, goal-directed behaviour, which relies on cognitive action planning and deliberate evaluation. Critically, goal-directed behavior allows for the flexible construction of value at each decision point through online planning procedures. This mechanism makes agents immediately sensitive to changes in outcome values, enabling them to dynamically adapt their actions to novel or changing circumstances [4]. Studies in network neuroscience have identified dynamic network reconfiguration, characterized by the flexible recruitment and integration of neural circuits, as a fundamental mechanism supporting this adaptability across cognitive[5] and motor domains[6].
Computational models within the predictive coding framework have recently provided insights into the neural mechanisms underlying dynamic adaptation. These models suggest that adaptation can be understood as a form of Bayesian inference, where the brain continuously infers latent contextual states to guide behavior [7]. A prominent example is the COIN model[8, 9], which posits that adaptation arises through two complementary processes: proper learning, involving the creation and updating of memories, and apparent learning, which dynamically adjusts the expression of pre-existing memories based on inferred context. While these frameworks capture key phenomena of adaptive behavior, such as context-dependent single-trial learning and spontaneous recovery, a significant gap remains between these theoretical models and the neural substrates that implement these computations in the brain.
@@ Introduction
In machine learning, the capacity of a model to adapt its behavior, without weight updates, is referred to as in-context learning (ICL). Initially, ICL was observed in architectures tailored for few-shot learning [10] or even zero-shot learning [11]. However, the game changed when it was observed that ICL emerges naturally in large-scale transformers [12, 13]. These architectures leverage self-attention, a mechanism that dynamically changes the way past inputs are attended, in order to shape the current response. Their exceptional capability to adapt to contextual information allowed transformer-based architectures to achieve state-of-the-art performances in many domains, such as natural language processing and image analysis[14, 15, 16]. While there are several intriguing explanations [17, 18] for the emergence of ICL in transformers, a more formal understanding was pioneered by Von Oswald and collaborators [19], who demonstrated that a linear self-attention layer can implement gradient descent steps within its forward pass. This intuition was empirically validated by analyzing trained transformers and was formally proven [20].
Despite their success, transformer architectures face notable shortcomings, particularly in their memory requirements, which scale quadratically with sequence length. These requirements make those architectures largely biologically implausible. Moreover, memory demands constrain their scalability and restrict their applicability to tasks involving long sequences, such as full-length document analysis or video understanding. To overcome these challenges, several efficient deep linear RNN architectures have been proposed [21, 22], demonstrating substantial performance gains over transformers, especially in long-sequence tasks.
However, these deep learning architectures rely on several mechanisms that do not have an obvious correspondence in biological neural systems (such as the attention mechanism itself), leaving an unresolved gap in relating them to the in-context learning abilities exhibited by biological networks. This work bridges this gap by introducing a constructive method to induce in-context learning in biologically plausible neural networks, demonstrating its applicability in a wide range of scenarios.
@@
Building on this premise, our investigation explores how in-context learning, as demonstrated in deep learning architectures, might also be realized in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We propose that in-context learning can be facilitated by input segregation and dendritic amplification, two features widely observed in biological networks. We further argue that these mechanisms provide a biologically plausible foundation for implementing a process analogous to the attention mechanism in transformers. Recent findings on dendritic computational properties [23] and on the complexity of pyramidal neurons dynamics [24] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [25, 26, 27, 28]. It has been proposed that segregation of dendritic input [26] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [28] would support backpropagation in biological neurons. Recently, it has been suggested [29] that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network. Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of temporal trajectories and reinforcement learning, involving non trivial input-output temporal relations. This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
Main / Framework description
Dynamical adaptation replaces synaptic plasticity: A popular example
Consider a task inspired by the Wisconsin card-sorting test. Each card in a deck is defined by two features: color and shape. In each iteration, the agent receives a cue card and two guide cards—one matching the cue by color and the other by shape. The agent must select the correct guide card to receive a positive reward $R(t)=1$. Initially, the correct choice is based on color matching. However, after 10 trials, the classification rule changes, rewarding shape matching instead. A successful agent must quickly adapt its decision-making to these contextual shifts.
We consider a simple neural network that predicts reward probability for each (cue card, guide card) pair. Two populations, $z_{1}$ and $z_{2}$, encode similarity based on color and shape, respectively. A third population, $y=\phi(w_{1}z_{1}+w_{2}z_{2})$, integrates this information to predict the reward probability. Here $\phi(x)=\text{sigmoid}(x-\text{bias})$ represents the neural transfer function. The weights $w_{1}$ and $w_{2}$ combine the features to determine the prediction, and the guide card with the highest predicted reward probability is selected. The agent dynamically updates the model based on received rewards to adapt to changing contexts.
Traditionally, weights $w_{1}$ and $w_{2}$ are considered synaptic weights, and learning is attributed to synaptic plasticity. Alternatively, $w_{1}$ and $w_{2}$ can be reinterpreted as the activity of two additional populations acting as gain modulators In our model, we refer to them as "virtual weights," as they replace traditional physical synaptic weights, effectively virtualizing the learning process. These modulatory populations dynamically reconfigure the responses of the network, enabling fast, flexible adaptations without the need for synaptic changes. One of the key advantages of this virtualization is that it allows for learning rules that are not constrained by the strict locality requirements inherent to synaptic plasticity.
To support this point, in Fig. 1 we compare synaptic learning using the gradient-based, delta rule $\Delta\mathbf{w}=\eta e\cdot\mathbf{z}$ with a non-local learning rule that force constraints $w_{1}(t)+w_{2}(t)=const$. This results in the update rule $\Delta w_{j}=\eta(-1)^{j}e\cdot(z_{2}-z_{1})$. Here $e(t)=(R(t)-y(t))$ represents the reward prediction error for the last choice and $\eta$ is the learning rate, controlling the speed of adaptation. As the latter learning rule features non-local terms, it can easily be implemented by a network that dynamically adjusts the activities of two modulatory populations of virtual weights $w_{1},w_{2}$ without relying on synaptic updates (see Fig. 1 C). In Fig. 1 we compare the performance of an agent that learns with synaptic plasticity with an agent that performs dynamic adaptation. As we can see in panel 1 F, decreasing the timescales of the learning process, the network with dynamic adaptation achieves maximum reward after seeing only one in-context example, while the performance of the network with synaptic plasticity plateaus at near chance level. This is supported by the psychometric curve shown in Fig. 1 G. Interestingly, the comparison between the two curves in Fig. 1 G closely mirrors findings from a recent study [30]. In that study, some human participants displayed the ability to adjust their behavior after just one example, while others, with impaired cognitive abilities, had to relearn the previous classification rule through trial and error from scratch. In this work, we propose biologically plausible framework for understanding rapid behavioral adaptations in dynamic environments, bridging the gap between biological and artificial intelligence.
Figure 1: Adapting behavior in a context-dependent classification task. A. Overview of the task structure: At each iteration, the agent is presented with a cue card and two guide cards, one matching the cue by color and the other by shape. The agent must select the correct guide card to receive a positive reward. The classification rule switches after 20 trials. B. Schematic representation of the proposed network architecture: synaptic plasticity of physical weights is replaced by dynamic adaptation of virtual weights C. that modulate the network response. D. Example temporal trajectory for the plastic network. Two in-context errors are needed to learn the correct rule. E. Example temporal trajectory for the network with dynamic adaptation: one in-context errors are needed to learn the correct rule. F. Average reward after 1 in-context example as a function of the learning rate $\eta$. Synaptic plasticity with a local delta rule does not go beyond the chance level. Dynamic adaptation with a non-local learning rule can achieve a perfect score. The curve is obtained by averaging over 10 independent sessions of 2000 iterations each. The shaded region represents the standard deviation confidence interval. G. Psychometric curve comparing average reward as a function of in-context examples. At $-1$ we report the average reward before the rule switch. The curve is obtained by averaging over 10 independent sessions of 2000 iterations each.
Framework Formalization
Consider a general learning scenario, where an agent is engaged in an environment $\alpha$ from a family of tasks $\mathcal{A}$. The agent receives temporal input $\mathbf{x}(t)$ and it has to compute a response $y(t)$, receiving feedback $f$ from the environment in return. This can be an error, reward, or the target behavior.
We further assume that a low-level network extracts features $\mathbf{z}(t)$ relevant to the task family. The agent then adapts its response to the specific environment by learning readout parameters $\mathbf{w}$, which determine how to recombine these features to produce the correct response, $y(t)=Y(\mathbf{z}(t),\mathbf{w})$. Here $Y$ is a generic function that depends on the parameters $\mathbf{w}$. These parameters are then adjusted at each time step in function of an internally computed error signal $e$. The overall learning process can then be generically rewritten as the following dynamical system:
$\begin{cases}\tau_{e}\,\dot{\mathbf{e}}&=E(\mathbf{e},\mathbf{y},\mathbf{f}),\\ \tau_{y}\,\mathbf{y}&=Y(\mathbf{y},\mathbf{w},\mathbf{z}),\\ \tau_{w}\,\dot{\mathbf{w}}&=W(\mathbf{w},\mathbf{z},\mathbf{e}).\end{cases}$ (1)
The specific instantiation of the functions $W$ and $E$ depends on the learning algorithm employed and can encompass both supervised and reinforcement learning settings. Refer to the Methods section for their precise definitions. In the classic view, the parameters $\mathbf{w}$ are interpreted as synaptic weights of a neuronal network, and the functions $E$ and $W$ describe synaptic plasticity mechanisms. However, this imposes stringent locality requirements, as all information must be present at the synapse.
Rather than interpreting $\mathbf{w}$ as synaptic weights, we view them as the dynamic activity of an auxiliary network that modulates the response $Y$ in real time based on feedback and past interactions. We refer to these dynamically adapting state variables as "virtual" weights. We refer to $\mathbf{w}$ as “virtual weights” because: - Weights: they replace the physical weights, and they have the same role as the physical weights of parametrizing the map from the features $\mathbf{z}(t)$ to the output(s). - Virtual: they are technically not weights, since the are encoded in the activity of the readout neurons of the network W, and their activity is provided as an input current to the Y network. Using this strategy we replace the computation done by the synapses (including learning) with operations performed by the neural activity. Dynamic adaptation offers key advantages over synaptic plasticity:
  • Fast Timescales: It naturally operates on faster timescales, making it well-suited for rapid behavioral changes
  • Non-Locality: Dynamic adaptation is not restricted by the stringent locality requirements of synaptic plasticity, allowing for richer forms of interaction and computation at the network level.
In this interpretation, the functions $E$ and $W$ in (1) arise from the dynamics of the auxiliary network of virtual weights. We show that such networks can dynamically implement standard gradient-based algorithms widely used in machine learning. While these algorithms are traditionally associated with synaptic plasticity, their reliance on non-local terms suggests that, in the brain, they can be more naturally realized at the circuit level. We demonstrate that such circuits capable of dynamic learning can simply arise by tuning the static readout weights of random RNNs with multiplicative interactions, requiring only a limited number of training trajectories from the original learning algorithm applied to a subset of the task family. The construction of our networks follows two distinct phases: a developmental or evolutionary phase, and a subsequent phase where the network leverages this structure for dynamical learning.
  • Phase 1: Development of the network structure. In this phase, using an Algorithmic Distillation protocol[31], we collect learning trajectories of the form $\{(\mathbf{z}_{\alpha}(t),\mathbf{w}_{\alpha}(t),\mathbf{e}_{\alpha}(t),\mathbf{f}_{\alpha}(t))_{t=0}^{T}\|\alpha\in\mathcal{A}_{ID}\}$ by running the learning dynamics in (1) on a subset of tasks $\mathcal{A}_{ID}$ (In Distribution). We then train the static readout weights $\Theta$ of two random RNNs $Y_{\Theta_{y}}$ and $W_{\Theta_{w}}$ to replicate, respectively, the functions $Y$ and $W$ on the training data.
  • Phase 2: The whole architecture is tested (dynamical learning). The architecture is tested on a new, previously unseen task $\alpha\in\mathcal{A}_{OOD}=\mathcal{A}\setminus\mathcal{A}_{ID}$ (Out-of-Distribution). During this phase, all synaptic weights remain fixed. Instead of relying on synaptic plasticity, the virtual weights evolve dynamically according to: $\mathbf{w}(t+dt)=\mathbf{w}(t)+W_{\Theta_{w}}(t)$. The updated output of the network is given by $y(t)=Y_{\Theta_{y}}(\{\mathbf{w}(t),\mathbf{z}(t)\})$. Importantly, we remark that $\mathbf{w}(t)$ no longer represents physical synaptic weights but rather contextual signals encoded in the activity of neurons, acting as input currents to the $Y_{\Theta_{y}}$ network.
The role of gain-modulation in adaptive behavior
Recent research has highlighted a growing interest in understanding modulatory mechanisms and their role in neural computation. It has been observed that gain modulation allows neurons to adjust their response to inputs based on context [24]. This dynamic regulation of neuronal sensitivity plays a critical role in adaptive behavior and cognitive flexibility. Moreover, it has been implicated in various functions, including feedback control [32, 33]) and attention, where it allows for the selective amplification of relevant information. A recent study [34] identified gain modulation as a key neural mechanism underlying hierarchical spatiotemporal prediction and sequence learning, demonstrating that such models can replicate various cortical processing features, including temporal abstraction and space-time receptive fields in the primary visual cortex. Furthermore, gain modulation is closely tied to meta-learning and multitask learning [35, 36], enabling neural networks to efficiently generalize across diverse tasks by leveraging contextual cues.
Within biological recurrent networks, dendritic segregation provides a structural basis for gain modulation. By compartmentalizing synaptic inputs, neurons can effectively implement non-linear interactions between different input streams. This is reminiscent of In-Context Learning (ICL) in machine learning, where multiplicative interactions are crucial for enabling models to leverage previously learned information in new contexts. Although we focus on gain modulation mediated by apical dendrites, other mechanisms are also known. For instance, neuro-glia interactions at the synapse have been proposed as a substrate for working memory, with astrocytes playing a crucial role in modulating neural activity and maintaining memory through glia–synapse interactions [37]. This gain modulation mechanism could complement other adaptive processes, such as subtle, correlated changes in connectivity [33].
Our work explores the role of gain modulation in recurrent networks, employing a simplified model where apical currents act multiplicatively on basal currents within a transfer function. A key parameter, $\gamma$, controls the strength of these non-linear interactions (see Methods for further details). By modulating the gain of neuronal responses, these networks exhibit enhanced adaptability and in-context learning capabilities.
Results